# Short geodesics and multiplicities of eigenvalues of hyperbolic surfaces

### v1

@@ 4.3.2. Estimate on $Y_{2}$
If $\ell<\epsilon/2$, then since $\ell<1$,
(38) |$\absolutevalue{f(p)-f(q)}^{2}$ $\leq\norm{\nabla f}_{L^{\infty}(B(p,\epsilon/2))}^{2}\cdot\ell^{2}\leq\frac{16}{9}\tilde{c}(\epsilon)^{2}\cdot\lambda\cdot\ell.$
If $\ell(\alpha)\geq\epsilon/2$, then for each integer $n\in\{0,\cdots,\lfloor 2\ell/\epsilon\rfloor\}$, set $\alpha_{n}=\alpha(n\epsilon/2)$. We have
(39) |$\absolutevalue{f(p)-f(q)}$ | $\leq\sum_{n=0}^{\lfloor 2\ell/\epsilon\rfloor-1}\absolutevalue{f(\alpha_{n})-f(\alpha_{n+1})}+\absolutevalue{f(\alpha_{\lfloor 2\ell/\epsilon\rfloor})-f(q)}$ |
$\leq\sum_{n=0}^{\lfloor 2\ell/\epsilon\rfloor}\norm{\nabla f}_{L^{\infty}(B(\alpha_{n},\epsilon/2))}\cdot\frac{\epsilon}{2}$
$\leq\frac{2}{3}\tilde{c}(\epsilon)\sum_{n=0}^{\lfloor 2\ell/\epsilon\rfloor}\norm{\mathop{}\!\mathrm{d}f}_{L^{2}(B(\alpha_{n},\epsilon))}\cdot\epsilon$
$\leq\frac{2}{3}\tilde{c}(\epsilon)\bigg{(}\sum_{n=0}^{\lfloor 2\ell/\epsilon\rfloor}\norm{\mathop{}\!\mathrm{d}f}^{2}_{L^{2}(B(\alpha_{n},\epsilon))}\bigg{)}^{1/2}\bigg{(}\sum_{n=0}^{\lfloor 2\ell/\epsilon\rfloor}\epsilon^{2}\bigg{)}^{1/2}.$
By Lemma 5.1, $B(\alpha_{n},\epsilon)\subset\mathcal{A}_{k}\cup\mathcal{S}_{k}$, therefore, if $B(\alpha_{n_{1}},\epsilon)\cap B(\alpha_{n_{2}},\epsilon)\neq\emptyset$, then |$\absolutevalue{n_{1}-n_{2}}<2$. Otherwise, we can find a shorter geodesic segment joining $p$ and $q$ that does not intersect with any $\gamma\in\mathcal{N}_{\epsilon}(X_{g})$, which leads to a contradiction. It then follows that
(40) $\sum_{n=0}^{\lfloor 2\ell/\epsilon\rfloor}\norm{\mathop{}\!\mathrm{d}f}^{2}_{L^{2}(B(\alpha_{n},\epsilon))}\leq 2\norm{\mathop{}\!\mathrm{d}f}_{L^{2}(X_{g})}^{2}=2\lambda.$
@@
(41) $\sum_{n=0}^{\lfloor 2\ell/\epsilon\rfloor}\epsilon^{2}\leq\sum_{n=0}^{\lfloor 2\ell/\epsilon\rfloor}\epsilon=(\lfloor 2\ell/\epsilon\rfloor+1)\epsilon\leq 4\ell,$
@@
(42) |$\absolutevalue{f(p)-f(q)}^{2}\leq\frac{32}{9}\tilde{c}(\epsilon)^{2}\cdot\lambda\cdot\ell.$
The lemma then follows from (38) and (42) by setting $c(\epsilon)=\frac{32}{9}\tilde{c}(\epsilon)^{2}$. ∎
@@
We consider the thick-thin decomposition of $X_{g}$ defined by (36). Inspired by the recent work [GLN24] of Gross-Lachman-Nachmias, for any $j\geq 0$, we also set
@@
$B_{x,R}:=\left\{z\in\mathcal{A}_{k}\cup\mathcal{S}_{k}:\ \begin{gathered}\text{there exists a geodesic segment of length $\leq R$ joining}\\ \text{$x$ and $z$ that does not intersect with any $\gamma\in\mathcal{N}_{\epsilon}(X_{g})$}\end{gathered}\right\}.$
@@
$\frac{\mu_{j-1}}{4}\leq\frac{f_{x}(x)^{2}}{4}\leq\absolutevalue{f_{x}(x)-f_{x}(z_{0})}^{2}\leq c_{2}\cdot R_{j}\cdot\lambda,$
@@
$\operatorname{Diam}_{c}(\mathcal{A}_{k}):=\sup_{x,z\in\mathcal{A}_{k}}\left\{\operatorname{\ell}(\alpha):\ \begin{gathered}\text{$\alpha$ is a shortest geodesic segment joining $x$ and $z$}\\ \text{that does not intersect with any $\gamma\in\mathcal{N}_{\epsilon}(X_{g})$}\end{gathered}\right\}.$
@@
and the geodesic balls $B(\alpha(nc_{1}/2),c_{1}/4)$, $11\leq n\leq\lfloor R/c_{1}\rfloor-$ 1, are pairwise disjoint. Moreover, by Lemma 5.1 we have that for all $10\leq n\leq\lfloor R/c_{1}\rfloor-$ 1,
@@
[…]
@@
$\leq 3C\sqrt{\lambda}\cdot\operatorname{Area}(X_{g})+24I_{\epsilon}.$
The proof is then finished by Gauss-Bonnet. ∎

### v2

@@ 4.3.2. Estimate on $Y_{2}$
If $\ell<\epsilon/4$, then since $\ell<1$,
[…]
If $\ell\geq\epsilon/4$, then for each integer $n\in\{0,\cdots,\lfloor 4\ell/\epsilon\rfloor\}$, set $\alpha_{n}=\alpha(n\epsilon/4)$. We have
[…]
$\leq\sum_{n=0}^{\lfloor 4\ell/\epsilon\rfloor}\norm{\nabla f}_{L^{\infty}(B(\alpha_{n},\epsilon/4))}\cdot\frac{\epsilon}{4}$
$\leq\frac{1}{3}\tilde{c}(\epsilon)\sum_{n=0}^{\lfloor 4\ell/\epsilon\rfloor}\norm{\mathop{}\!\mathrm{d}f}_{L^{2}(B(\alpha_{n},\epsilon/2))}\cdot\epsilon$
$\leq\frac{1}{3}\tilde{c}(\epsilon)\bigg(\sum_{n=0}^{\lfloor 4\ell/\epsilon\rfloor}\norm{\mathop{}\!\mathrm{d}f}^{2}_{L^{2}(B(\alpha_{n},\epsilon/2))}\bigg)^{1/2}\bigg(\sum_{n=0}^{\lfloor 4\ell/\epsilon\rfloor}\epsilon^{2}\bigg)^{1/2}.$
[…]
(40) $\sum_{n=0}^{\lfloor 4\ell/\epsilon\rfloor}\norm{\mathop{}\!\mathrm{d}f}^{2}_{L^{2}(B(\alpha_{n},\epsilon))}\leq 2\norm{\mathop{}\!\mathrm{d}f}_{L^{2}(X_{g})}^{2}=2\lambda.$
@@
(41) $\sum_{n=0}^{\lfloor 4\ell/\epsilon\rfloor}\epsilon^{2}\leq\sum_{n=0}^{\lfloor 4\ell/\epsilon\rfloor}\epsilon=(\lfloor 4\ell/\epsilon\rfloor+1)\epsilon\leq 8\ell,$
@@
[…]
The lemma then follows from (38) and (42) by setting $c(\epsilon)=\frac{16}{9}\tilde{c}(\epsilon)^{2}$. ∎
@@
We consider the thick-thin decomposition of $X_{g}$ defined by (36). Inspired by the recent work [GLN2024] of Gross–Lachman–Nachmias, for any $j\geq 0$, we also set
@@
[…]
@@
[…]
@@
[…]
@@
and the geodesic balls $B(\alpha(nc_{1}/2),c_{1}/4)$, $1\leq n\leq\lfloor R/c_{1}\rfloor-1$, are pairwise disjoint. Moreover, by Lemma 5.1 we have that for all $0\leq n\leq\lfloor R/c_{1}\rfloor-1$,
@@
$\int_{\mathcal{A}_{k}}f_{x}(x)^{2}\leq\max\Big\{C\sqrt{\lambda}\cdot\operatorname{Area}(\mathcal{A}_{k}\cup\mathcal{S}_{k}),\ 8\Big\}.$
@@
$\leq 3\sum_{k=1}^{I_{\epsilon}}\max\Big\{C\sqrt{\lambda}\cdot\operatorname{Area}(\mathcal{A}_{k}\cup\mathcal{S}_{k}),8\Big\}$
@@
The proof is then finished by Gauss–Bonnet. ∎